if an isosceles triangle has congruent sides of 15 cm and the base is 20 what would the hieght be
Answer by swincher4391(1107)

Imagine you have two scalene triangles put together with leg 10 and hypotenuse 15
Then, 10^2 + b^2 = 15^2 by the Pythagorean theorem. 100 + b^2 = 225
b^2 = 125
b = sqrt(125) = 5*sqrt(5)
I hope that this helped!
If this was not detailed enough please e-mail me at onlyyourmathtutor@yahoo.com.